OpenStudy (anonymous):

Surface area of a cylinder with a radius of 4 and height of 5 (must be in terms of pi)

OpenStudy (anonymous):

do u know its formula

OpenStudy (anonymous):

\[2 \pi rh+2 \pi r^2\] just plug in the values r=4, h=5 u'll get the surface area.

OpenStudy (anonymous):

yeah but I get to 2pi16+2pi20 and don't know where to go from there

OpenStudy (anonymous):

Like, what do I multiply to get the final answer?

OpenStudy (anonymous):

@kmillz73

OpenStudy (anonymous):

yes? @sheelo_mughal all you said was my name.